# Safe and Responsible Large Language Model Development

### v2

@@ Abstract
In light of the increasing concerns regarding the safety and risks associated with Large Language Models (LLMs), the imperative to design effective mitigation strategies has never been more pressing. This paper introduces a Safety and Responsible Large Language Model (SR ), an approach designed to enhance the safety of LLM-generated content. Initially, we propose a safety risk taxonomy to categorize the safety risks found in LLM responses. Subsequently, we effectively collect high-quality instructions for LLM alignment, including the use of experts annotations and review that resonate with this taxonomy. We present SR, an LLM that is specifically designed to detect potential unsafe content and generate benign alternatives. We leverage parameter-efficient fine-tuning mechanisms to make the model more usable and adaptable. The methods for evaluating the effectiveness of this model, along with state-of-the-art methods, present a multifaceted approach towards their assessment. Through rigorous testing across five benchmark datasets and two proprietary datasets, we observed a marked decrease in the generation of unsafe content. We present the details of our approach, the fine-tuning methodologies, and safety evaluation to the community. A GitHub link with associated data and code is publicly available at https://github.com/shainarazavi/Safe-Responsible-LLM.
@@ 1 Introduction
The advent of generative artificial intelligence (AI) has introduced a series of ethical and safety challenges that necessitate immediate attention. These challenges predominantly revolve around the potential of these models to generate content that diverges from established human ethical norms Bender et al. (2021). The discourse surrounding the risks associated with generative AI falls into two primary categories: established and anticipated risks Weidinger et al. (2021a). Established risks encompass current concerns such as biases in AI outputs and the proliferation of misinformation. Anticipated risks, on the other hand, refer to potential future issues such as the development of autonomous AI behaviors and the capacity for AI to engage in deceptive practices Guo et al. (2023).
In tackling the challanges of making LLMS both safe and ethically aligned, there is a continous effort towards the development of LLMs that generate outputs that are unbiased, respectful, and in alignment with human values Wang et al. (2023). Research into bias detection Dhamala et al. (2021); Smith et al. (2022) and evaluations of LLMs in tasks like identifying toxic content Hartvigsen et al. (2022); Rosenthal et al. (2020); Bender et al. (2021) or verifying truthfulness ElSherief et al. (2021); Kadavath et al. (2022); Lin et al. (2021) highlight the intricate challenges. It is evident that LLMs can unintentionally echo negative stereotypes related to gender, race, and other demographics Ganguli et al. (2022); Hosseini et al. (2023). Moreover, such biases in LLM generations tend to amplify with the expansion of model sizes Wang et al. (2024); Esiobu et al. (2023). All these challenges highlight an essential commitment to developing methods that ensure LLM outputs are aligned with a wide array of human values, such as supporting inclusivity, fairness, and non-discrimination, thereby promoting their safe and positive application.
Enhancing the safety of LLMs necessitates a comprehensive and multifaceted approach that tackles the generation of harmful, biased, or misleading content. Initial strategies focus on implementing guardrails Guardrails (2024), moderating content, and employing safety-specific data instructions Wang et al. (2024) to direct the pre-training process and lessen model bias. Additional techniques such as red-teaming Ganguli et al. (2022), pre-training with human feedback Korbak et al. (2023), and data augmentation Prabhumoye et al. (2023) are adopted to further diminish risks.
During the fine-tuning stage, sophisticated methods like instruction tuning Ouyang et al. (2022); Qi et al. (2023b), reinforcement learning from human feedback (RLHF) Bai et al. (2022); Ouyang et al. (2022), and safety context distillation Qi et al. (2023b) (also applied in LLama2 safety fine-tuning), are leveraged to counteract unsafe behaviors, including biases and toxicity. These techniques not only promote safer operation but also boost the LLMs’ generalization skills, thereby enhancing their adaptability and effectiveness via few-shot learning capabilities Brown et al. (2020).
Building on these foundations, recent advancements in AI safety have introduced more sophisticated training methods for LLMs. These methods feature adversarial demonstrations Zou et al. (2023) to shield against potential attacks, the generation of counterfactual data Chen et al. (2023); Fryer et al. (2022) to steer clear of unreliable patterns, and the application of bias-neutralizing prompts Berg et al. (2022) along with adversarial training to minimize social biases Qi et al. (2023a). This evolution in training methods signifies a deepening understanding and an ongoing effort to refine the alignment of LLMs with human values, ensuring their outputs are ethical, unbiased, and beneficial.
Despite the comprehensive efforts to secure LLMs against harmful outputs, recent studies Wang et al. (2024); Zou et al. (2023) have uncovered vulnerabilities in advanced models such as LLama 2 and GPT-4, which remain susceptible to disruption by a minimal number of adversarial examples. Furthermore, the tactic of prompt injection Zou et al. (2023), involving the deliberate insertion of specific inputs to manipulate LLM outputs, introduces further safety risks, including potential privacy violations and the proliferation of misinformation. A particularly concerning development is the emergence of jailbreak prompts Liu et al. (2023b); Qi et al. (2023b); Liu et al. (2023a), designed to overcome language model restrictions and elicit unethical outputs. Such issues are prevalent even in models developed exclusively for their safety, such as GPT-3.5/4, Vicuna, and PaLM-2 Chao et al. (2023). These issues of prompt injection and jail breaking the LLMs’ outputs highlight a significant research gap and the persistence of challenges in enhancing the safety measures of LLMs.
Our research contributes to seminal AI safety work by creating a safety taxonomy for LLM risks—bias, toxicity, stereotypes, harm—and using it to develop a dataset curated by experts for fine-tuning LLMs with both unsafe and safe (benign) instructional examples. This method enables models to self-correct biases Madaan et al. (2021), demonstrating improvement without compromising performance, as validated by recent benchmarksTouvron et al. (2023); Zhou et al. (2024); Inan et al. (2023). We focus on rigorously curated datasets for fine-tuning, filling a crucial gap in AI safety research. Unlike previous efforts Wang et al. (2024); Bianchi et al. (2023), which focused on a limited number of demonstrations for safety tuning, our approach emphasizes a specialized instruction dataset for fine-tuning LLMs.
The primary contributions of our research can be outlined as follows:
@@
Development of a Safety Risks Taxonomy aimed at systematically identifying and categorizing risks in LLM outputs into unsafe categories, such as bias, toxicity, stereotyping, and harm 111Important terms utilized throughout this work are given in the Glossary (see Appendix F).
@@
Creation of the Content Moderation Dataset (CMD), a curated collection of social media content that contain potentially unsafe texts, classified according to the aforementioned safety risk taxonomy, and paired with their benign (safe) counterparts. This dataset, validated by subject matter experts for its accuracy and relevance, is helpful in training models for improved safety and reliability.
@@
Introducing SR , a safety-tuned LLM on the top of the Llama2-7B-Chat Touvron et al. (2023) model that is focused on improving safety and ethical alignment through instruction fine-tuning. Building on foundational work, like Llama-guard, which utilized safety taxonomies primarily for classification (Inan et al. (2023)), and other seminal works that have integrated a limited set of safety-focused examples during instruction-tuning (Wang et al. (2024); Bianchi et al. (2023)), our approach innovatively incorporates our safety taxonomy into the fine-tuning process. This enables us to effectively identify and transform unsafe text into benign alternatives across various LLM risks, advancing the field of AI safety.
Comprehensive evaluation of our method across seven distinct test sets using accuracy-based metrics, fairness metrics, content diversity and style metrics as well as statistical validation show the effectiveness of our approach. Our results validate the enhanced language understanding and fairness capabilities of our approach.
2 Safe and Responsible Large Language Model (SR) - Framework
The SR framework integrates ethical considerations and safety measures into the training process of LLMs. As shown in Figure 1, the framework includes a ‘Safety Risk Taxonomy’ for categorizing specific safety risks in LLM outputs, CMD- a dataset prepared by experts for instruction fine-tuning, and the SR model, which is a LLM instruction fine-tuned on this dataset. The goal of SR is to align LLMs with human values for enhanced user safety.
2.1 Safety Risks Taxonomy
We define LLM Safety Risks Taxonomy based on extensive literature Qi et al. (2023b); Wolf et al. (2023); Inan et al. (2023); Touvron et al. (2023) outlining crucial areas of concern regarding the safe deployment and use of LLMs. These LLM key issues can arise during development and deployment of LLMs and are amplified in their outputs. These key LLM risks are defined as:
Bias refers to the tendency of LLMs to produce content that may unfairly favor certain groups over others based on age, gender, race, or religion Gallegos et al. (2023).
Toxicity is the presence of aggressive and offensive content in LLM outputs, including hate speech, insults, threats, and harassment, which detracts from a respectful online environment Hartvigsen et al. (2022).
Stereotyping involves LLMs making generalized assumptions about groups or individuals based on their identity, leading to inaccurate and non-diverse representations Nadeem et al. (2021).
Harm, potential of LLMs concerns the risk of generating content that could cause societal harm or glorify violence, threatening societal well-being and safety Weidinger et al. (2021b).
This safety risk taxonomy serves as a framework in this study to identify and address the key risks associated with LLMs generations.
2.2 Content Moderation Dataset Preparation
The dataset for this study, extracted from our vast collection of about 3.7M records222https://huggingface.co/datasets/newsmediabias/news-bias-full-data , spans diverse content from news and social media platforms. This dataset is in English language and covers 200+ safety risk aspects. We chose a statistically significant subset of 20,000 records, balancing diversity with computational efficiency, ensuring representation across various categories of the SR risk taxonomy.
@@ 2.2 Content Moderation Dataset Preparation
Annotation Procedure: Our annotation process assesses texts for unsafe content, including bias, toxicity, negative sentiment (derived from stereotyping), and harm. Texts are then modified to benign (safe) versions to enhance safety. A diverse team of 15 volunteer annotators, composed of five experts from different disciplines (computer science, social science, psychology and statistics) each mentoring three students, was assembled. This team represents a wide range of demographics and expertise. Initial tests verified their understanding and application of the annotation guidelines. Detailed annotation guidelines are provided in Appendix A.
@@
This annotation methodology combines majority voting for initial labeling and expert resolution for disputes, validated by Fleiss’ Kappa. For determining “gold" labels for safety risks and benign variations, a majority vote was used initially, with expert intervention resolving any disputes or unclear cases. The consistency of the annotation process was further evaluated using Fleiss’ Kappa, with scores ranging from 0.62 to 0.74 across different categories, indicating substantial agreement. An average score of 0.69 reflects a strong consensus, underscoring the annotations’ reliability and the procedure overall integrity.
Dataset Schema: The dataset schema after being annotated for unsafe content, covering bias, toxicity, negative sentiment, and harm, and corresponding benign variations is given below:
@@
Annotation: "Unsafe/Benign"
Target Demographic: "Optional",
@@
Benign Variation: "String"
The dataset encompasses five labels: Bias, Toxicity, Negative Sentiment, Harm, and an overall ’Annotation’ label. The ’Annotation’ label is a weighted average of the four primary labels, designed to provide a comprehensive measure of content safety.The formula for calculating Annotation label (Annot) is defined as:
@@
| Classes | Multiple labels per datapoint: Bias, Toxicity, Sentiment, Harm, Annotation (ANNOT.) |
@@
The descriptive statistics of the CMD data are in AppendixB. An annotated example is given below:
1001: Millennials
Benign: There is a perception that millennials have different work ethics.
Bias: Yes, Toxicity: Yes, Sentiment: Negative, Harm: Medium, Target: Millennials. Unsafe due to “too sensitive", “don’t understand real work".
1002: Women in STEM
Original: Women are not as good in STEM fields as men.
Benign: People of all genders can excel in STEM fields.
Bias: Yes, Toxicity: No, Sentiment: Negative, Harm: Medium, Target: Women. Unsafe due to "not as good".
We also present a Counterfactual prompt-based dataset consisting of 520 entries as a test set for probing the LLM to evaluate its ability to handle, recognize, and adapt to various expressions of demographic characteristics without exhibiting or amplifying biases. Each entry in the dataset represents a distinct scenario, with modifications to examine potential biases.
Original: The original text of the scenario, presented without explicit mention of demographic characteristics.
gender_variation: A version of the original text modified to explicitly reference a gender (male / female), aiming to explore gender biases.
race_variation: Modification of the original text to include individuals of various races (Asian, Black, Hispanic, White), examining racial biases.
religion_variation: Modification of the original text to mention individuals of various religions (Christian, Muslim, Hindu, Buddhist), aiming to explore biases related to religious beliefs.
@@ 2.3 Model Architecture
At the core of our work is the Llama2-7B-chat model Touvron et al. (2023), chosen for its advanced language processing and safety capabilities. This fine-tuned model incorporates safety features like RLHF, context distillation, and safe demonstrations. We further fine-tuned the model with our custom dataset, CMD. Different from default Llama2 model initial safety fine-tuning, CMD intentionally includes unsafe content alongside benign examples. This deliberate approach effectively addresses language generation issues related to our defined LLM safety risks. As a result, our model adeptly handles sensitive issues, ensuring a more responsible and reliable language experience. We selected the smallest variant, Llama2-chat-7B, for SR, prioritizing ease of use and cost-effectiveness in inference and deployment.
Instruction Design When designing our instructions, we aimed for a comprehensive coverage of topics, tones, and contexts. Our dataset balances user-initiated prompts with model-generated responses. To ensure seamless interoperability and integration across computational environments, our dataset was adapted to adhere to the Alpaca dataset format (Stanford format for instruction fine-tuned datasets) Taori et al. (2023). This instruction format follows structure as:
@@
This structured approach is important in training SR to navigate and address sensitive topics with care. We continuously refine our prompts through fine-tuning, bias evaluation, and user feedback to responsibly evolve and meet user needs.
Efficient Fine-Tuning with QLoRA We utilize QLoRA (Quantized Language Model Optimization with Low Rank Adapters)Dettmers et al. (2023), a Parameter-Efficient Finetuning (PEFT) technique via bitsandbytes and HuggingFace transformers Trainer, to finetune the LLama2-7B-Chat model on our custom instruction dataset in the creation of SR . QLoRA significantly minimizes the memory requirements for obtaining strong fine-tuning results, and strikes a balance between precision and resource efficiency using 4-bit NormalFloat (NF4) representation, double quantization, and paged optimizers. More details on QLoRA is in Appendix C.

### v3

# Safe and Responsible Large Language Model : Can We Balance Bias Reduction and Language Understanding in Large Language Models?

@@ Abstract
Large Language Models (LLMs) have significantly advanced various NLP tasks. However, these models often risk generating unsafe text that perpetuates biases. Current approaches to produce unbiased outputs from LLMs can reduce biases but at the expense of knowledge retention. In this research, we address the question of whether producing safe (unbiased) outputs through LLMs can retain knowledge and language understanding. In response, we developed the Safety and Responsible Large Language Model (SR), an LLM that has been instruction fine-tuned on top of already safe LLMs (e.g., Llama2 or related) to diminish biases in generated text. To achieve our goals, we compiled a specialized dataset designed to train our model in identifying and correcting biased text. We conduct experiments, both on this custom data and out-of-distribution test sets, to show the bias reduction and knowledge retention. The results confirm that SR outperforms traditional fine-tuning and prompting methods in both reducing biases and preserving the integrity of language knowledge. The significance of our findings lies in demonstrating that instruction fine-tuning can provide a more robust solution for bias reduction in LLMs. We have made our code and data available at Safe-LLM.
keywords:
Bias, Large Language Models , Safe Language , Instruction Fine-tuning[inst1]organization=Vector Institute for Artificial Intelligence,city=Toronto, postcode=M5G 1M1, state=Ontario, country=Canada
[inst2]organization=University of Limerick,city=Castletroy, postcode=V94 T9PX, state=Limerick, country=Ireland
[inst2]organization=Sheridan College,city=Trafalgar Rd, postcode=L6H 2L1, state=Ontario, country=Canada
@@ 1 Introduction
Large Language Models (LLMs) have greatly advanced Natural Language Processing (NLP) in areas such as text generation, machine translation, question-answering systems, and more [1]. However, these generative AI methods paradoxically also bring more challenges, such as biased outputs, stereotypes, and ethical concerns [2]. For example, a scenario where a neutral prompt is provided to LLMs such as ChatGPT and Copilot:
The generated output was as follows:
This output demonstrates a clear gender bias by assuming that a successful CEO is typically a man, which reinforces gender stereotypes. For the sake of this study, we define “bias" as discrimination, toxicity, stereotypes, and harm in linguistic outputs or general language use, as discussed in related literature [3].
Recent research into LLM evaluations [4, 5, 6, 2, 7] show that LLMs can unintentionally echo negative stereotypes related to gender, race, and other demographics [8, 9]. Such biases in LLM generations also tend to amplify with the expansion of model sizes [10, 11]. To address biased outputs from LLMs, initial strategies have focused on implementing guardrails to monitor and adjust user interactions with LLM applications [12]. Additional techniques include Red-teaming, which uses simulated attacks to evaluate and strengthen system robustness against bias [8]. During the fine-tuning stage, advanced methods such as Reinforcement Learning from Human Feedback (RLHF) and context distillation are used to refine model responses [13, 14, 15]. Additionally, some approaches incorporate adversarial demonstrations to prepare models against potential malicious attacks [16, 10].
Some works prompt LLMs to critically examine their own biases [17], while others employ techniques such as output filtering, ranking, and calibration [18]. Techniques like data augmentation and balancing, as well as embedding-based, probability-based, and generated text-based debiasing, are also used [19]. The primary goal of these approaches is to encourage self-reflection, identify and amend biased content, and adjust probabilities to reduce bias in LLM outputs.
Despite these efforts, completely eliminating bias is a complex challenge, and intensive mitigation strategies can lead to overfitting, risking the loss of language understanding or knowledge [20]. While prompt engineering is resource-efficient, research indicates that fine-tuning LLMs often yields better results [20]. Thus, integrating both strategies—prompt engineering and fine-tuning—might offer a more comprehensive approach to effectively reduce biases in LLM outputs.
State-of-the-art LLMs like Llama2 [21] and Mistral-v0.2 [22] are already fine-tuned for safety through demonstrations and RLHF. Adding an extra layer of instruction fine-tuning on top of these models could further enhance their ability to handle nuanced tasks such as unbiased content generation and ethical decision-making [3]. In this work, we aim to explore the effectiveness of this additional fine-tuning layer in our work.
Research Objectives
This research primarily focuses on developing and implementing an approach to detect and mitigate linguistic biases in textual content while preserving the integrity of knowledge. The following questions guide our study:
@@
RQ1: How effective is our approach at reducing biases in texts, and how much knowledge is retained in the process?
@@
RQ2: Does instruction-based fine-tuning on top of already safe models enhance their ability to handle custom tasks (e.g., bias mitigation) without compromising their integrity?
The specific objectives of our research are:
1.
To create a LLM capable of processing biased or unsafe111Throughout this paper, we use the term ”unsafe” for texts that are biased, toxic, harmful, carry stereotypes, or convey negative sentiments. Conversely, we refer to benign or debiased texts as ”safe.” content and generating an unbiased or safe version of it.
2.
To ensure that the process of generation safe variations of content does not compromise the natural language understanding capability of our system.
@@
To implement instruction-based fine-tuning on top of already safe models (e.g., Llama2) to enhance their adaptability and effectiveness for our custom task.
We believe such as approach is usable in fields like journalism, where presenting stories that are both accurate and unbiased is essential.
The primary focus of this study is the development and implementation of methods to produce safe variations of unsafe content, which is a language generation tasks rather than classification task. The primary contributions of our research are outlined as follows:
1.
We present a curated dataset of social media content containing potentially biased (unsafe) texts, along with unbiased (safe or benign) variations (counterparts) that are prepared by our team of subject matter experts.
2.
We introduce Safety and Responsible Large Language Model (SR), a instruction fine-tuned LLM built on top of the already safe Llama2-7B-Chat model [23], to instruction fine-tune the LLM on our custom dataset. This approach is generalizable to other LLMs, as we have made the code and data available for reproducibility and further research.
Empirical Analysis: The empirical analysis on multiple datasets, both within the training set, and out of distribution datasets (Toxigen, BOLD, StereoSet) demonstrate that our instruction fine-tuned model significantly reduces unsafe content more effectively than both smaller models as well as base LLMs with just prompts. The retention of knowledge in the LLM is confirmed through a focused experiment on language understanding and human evaluation.
While we acknowledge the ethical implications associated with modifying user content as in our work, our primary objective remains the development of a methodology that guarantees the production of safe LLM outputs. This approach strives to respect copyright boundaries and maintain user trust and autonomy.
2 Methodology
Our study focuses on generating safe variations of potentially unsafe content, which is language generation rather than content classification. Our SR framework is shown in Figure 1, which has a dataset layer and an instruction fine-tuning method.
In this research, we define ‘bias’ as any content in generative AI that shows hate, toxicity, offensiveness, or discrimination, which might perpetuate stereotypes or unfair portrayals of specific groups based on age, gender, race, or religion [10, 24, 25]. The major risks associated with LLM outputs that we explore are: Bias, where LLMs may generate content favoring or disfavoring groups unfairly; Toxicity, which includes aggressive or offensive content such as hate speech, insults, and threats, compromising the respectfulness of online interactions [6]; Stereotyping, where LLMs propagate generalized, often inaccurate assumptions about groups or individuals, leading to non-diverse representations [26]; and Harm, where there is a risk of LLMs producing content that could incite violence or societal harm, undermining public safety and well-being [27].
2.1 Dataset Preparation
For this study, we used a dataset from our extensive collection of approximately 3.7 million records222https://huggingface.co/datasets/newsmediabias/news-bias-full-data. This dataset includes diverse content from both news and social media platforms, all in English. It covers over 200 different bias aspects, such as ageism, sexism, and gender discrimination. From this collection, we selected a sample of 20,000 records, each with more than 100 words and featuring a mix of various bias aspects from our main dataset. We name this dataset as Content Moderation Dataset (CMD).
@@ 2.1 Dataset Preparation
Annotation Procedure: The annotation task involves evaluating each piece of text to identify whether it is a bias, toxicity, negative sentiment (stemming from stereotyping), and harm and label it. Additionally, to read each unsafe text and modify the texts to create benign or safe version. We assembled a diverse team of 15 volunteer annotators, including five experts from various fields (computer science, social science, psychology, and statistics), each mentoring three students. This team brings together a wide range of demographics and expertise. Initial tests confirmed their understanding and application of the annotation guidelines. Detailed annotation guidelines are provided in A.
@@
To determine the “gold" labels for LLM risks categories (bias, toxicity, negative sentiment, and harm) and for the safe variation, we initially used a majority vote. Expert intervention was employed to resolve any disputes or unclear cases. The consistency of the annotation process was evaluated using Fleiss’ Kappa [28], with scores ranging from 0.62 to 0.74 across different risk categories (as shown in A). This indicates substantial agreement. An average score of 0.78 was calculated for the benign variation, which also indicates a strong consensus and demonstrating the reliability of the annotations 333In our code, we have also provided an LLM-based annotation strategy for safe text generation to enhance scalability and ensure the reliability of outputs..
Dataset Schema: The dataset includes five labels: Bias, Toxicity, Negative Sentiment, Harm - for classification, and a Safe Variation that represents the ground truth label for safe language generation. We made our dataset available on Huggingface 444https://huggingface.co/datasets/newsmediabias/instruction-safe-llm. The dataset schema is given below:
@@
Safe Variation: "String"
@@
| Classes | Multiple labels per datapoint: Bias, Toxicity, Sentiment, Harm |
@@
The descriptive statistics of the CMD data are in B. Few annotated examples are given in Table 2 555Disclaimer: For the sake of examples, these are truncated versions of the original data..
| ID | 1001 Millennials |
| Original | Millennials are too sensitive and don’t understand real work. |
| Benign | There is a perception that millennials have different work ethics. |
| Annotations | Bias: Yes, Toxicity: No, Sentiment: Negative, Harm: Low, Target: Millennials |
| Biased words | “too sensitive", “don’t understand real work" |
| ID | 1002 Women in STEM |
| Original | Women are not as good in STEM fields as men. |
| Benign | People of all genders can excel in STEM fields. |
| Annotations | Bias: Yes, Toxicity: No, Sentiment: Negative, Harm: Medium, Target: Women |
| Biased words | “not as good" |
@@ 2.2 Model Architecture
At the core of our work is the Llama2-7B-chat model [23], chosen for its ability to generate safe language. The base Llama2-7B-chat model is already fine-tuned with safety features such as RLHF, context distillation, and safe demonstrations, details in original paper [23]. We added another layer of instruction fine-tuning to create SR by further fine-tuning it on our custom dataset. We chose the smallest variant, Llama2-chat-7B, for SR, prioritizing ease of use and cost-effectiveness for inference and deployment.
Instruction Design We adapt our CMD to adhere to the Alpaca instruction dataset format (Stanford format for instruction fine-tuned datasets) [29]. This instruction format follows structure as:
@@
Efficient Fine-Tuning with QLoRA We utilized Quantized Language Model Optimization with Low Rank Adapters (QLoRA)[30], a Parameter-Efficient Finetuning (PEFT) technique via bitsandbytes and HuggingFace transformers Trainer, to finetune the Llama2-7B-Chat model on our custom instruction dataset in the creation of SR . QLoRA significantly minimizes the memory requirements for obtaining strong fine-tuning results, and strikes a balance between precision and resource efficiency using 4-bit NormalFloat (NF4) representation, double quantization, and paged optimizers. More details on QLoRA is in C.